OpenStudy (anonymous):

A local AM radio station broadcasts at a frequency of 751 KHz. Calculate the energy of the frequency at which they are broadcasting.

OpenStudy (anonymous):

To calculate the energy of a electromagnetic wave, we use the formulas \(E=\frac{hc}{\lambda}\) or \(E=h\nu\) \(h\) is the Planck's constant, \(6.626\times 10^{-34}\) J.s \(c\) is the speed of light, \(3\times 10^8\) m/s \(\lambda\) is the wavelength, in meter. \(\nu\) is the frequency, in hertz. Frequency was given so you should use the second formula.